# Inclusive Training Separation and Implicit Knowledge Interaction for Balanced Online Class-Incremental Learning

### v1

@@ 5.3 Ablation Studies
w/o DC $58.2$ $21.6$ $83.1$ $24.2$
w/ DC $58.8$ $15.1$ $80.7$ $35.2$
5.4 Effectiveness of Inclusive Training Separation
Although we have confirmed the effectiveness of dual classifiers (DC) in our ablation study shown in Table 3, we further conduct an in-depth analysis on the benefit of our inclusive training separation using DC. To this end, in Figure 5, we present the confusion matrix, when solely applying DC to the baseline method (i.e., vanilla ER with the NCM classifier) on Split CIFAR-10 with a buffer size of 1k, in comparison with those without using DC. In confusion matrices, the x-axis represents predictions made by the models and the y-axis denotes true class labels, where we can also get class-wise accuracies from the diagonal values. The results are obtained by making predictions against 10k test samples, 1k per class, after training five incremental tasks of Split CIFAR-10, which are (ship, dog), (frog, cat), (horse, automobile), (truck, bird), and (deer, airplane) in order.
Reducing Bias Towards New Tasks.  Our first observation from Figure 5 is that predictions for the newest classes, deer and airplane, are reduced when using DC, as shown in Figure 5(b) (2,404 predictions in total), compared to when DC is not used, as shown in Figure 5(a) (2,932 predictions in total). This reduction indicates that our training separation effectively mitigates bias towards new tasks. Furthermore, this approach enhances the precision of predictions for these new classes: the precision for airplane improves from $\frac{748}{1,351}=55.4\%$ to $\frac{568}{888}=64.0\%$, and for deer from $\frac{591}{1,581}=37.4\%$ to $\frac{570}{1,516}=37.6\%$.
Capturing Semantic Relationship.  Figure 5 also indicates that our inclusive separation effectively captures semantic relationships between classes, as expected in Figure 3. For example, predictions for semantically similar classes, such as automobile and truck or cat and dog, become more correlated in Figure 5(b) compared to Figure 5(a). This high level of correlation suggests that our inclusive separation scheme makes it more difficult for the model to differentiate between similar classes. However, this increased correlation is what we intend to achieve for better feature embeddings, as it helps reduce incorrect predictions for unrelated classes, thereby enhancing overall accuracy. This improvement is evident in Figure 5, where there is a notable reduction in incorrect predictions to unrelated classes, such as from cat to frog (from $205$ in Figure 5(a) to $118$ in Figure 5(b)), despite the increased misclassification between similar classes like cat and dog. Consequently, as observed from the diagonal values, the class-wise accuracy for cat increases from $24.2\%$ to $35.2\%$, and for dog, it rises from $29.9\%$ to $43.2\%$. Moreover, we present the average accuracy (AA) among 10 classes, the corresponding standard deviation (STD), the maximum and minimum class-wise accuracy in Table 4. The method fully employing DC not only maintains a better overall accuracy, but also exhibits more balanced performance across all classes, from the smaller STD, the smaller Max, and the larger Min values, compared to the version without DC. This shows the effectiveness of inclusive separation of DC in improving overall performance as well as achieving balanced outcomes.
6 Conclusion
In this work, we proposed a novel method for online class-incremental learning, named BOIL, to overcome the limitations of existing replay-based methods, which often favor either plasticity or stability. BOIL suggests employing dual classifiers for inclusive separation of learning new samples from replaying buffer samples, thereby mitigating bias toward new tasks yet effectively consolidating knowledge across different tasks. By introducing implicit knowledge exchange between the dual classifiers, BOIL facilitates both forward and backward transfer throughout the incremental learning process. Empirical results demonstrate that BOIL achieves the best classification performance by providing the most balanced learner for OCIL.

### v2

@@ Acknowledgements.
PCR $29.8\pm 0.7$ $36.3\pm 0.6$ $43.5\pm 0.5$
BISON(ours) $33.5\pm 0.9$ $40.4\pm 0.8$ $45.5\pm 1.3$
Figure A3. Average accuracy of our BISON on Split CIFAR-100 ($\mathcal{M}$ = 2k) with various sizes of buffer batch.
Table A4. Analysis of computational efficiency on Split CIFAR-10 ($\mathcal{M}$ = 1k) and Split CIFAR-100 ($\mathcal{M}$ = 5k) datasets.
Method Split CIFAR-10 ($\mathcal{M}$ = 1k) Split CIFAR-100 ($\mathcal{M}$ = 5k)
Num of Params. GPU Usage(Mb) Num of Params. GPU Usage(Mb)
ER 1094750 1787 1109240 1837
BISON 1096343 1805 1125143 1853
Impact of Size of Buffer Batch $|\mathcal{B}_{\mathcal{M}}|$ . In our experiments, buffer batch size $|\mathcal{B}_{\mathcal{M}}|$ is set to $10$, which is the same as the size of stream batch. As shown in Figure A3, we examine the impact of buffer batch size on classification performance. As the buffer batch size becomes larger, the performance gradually improves. However, when it is larger than $32$, the improvement is limited. Remarkably, when the model is trained in an online setting, both efficiency and performance need to be taken into consideration, and hence it is not always feasible to set this value too large.
Computational Efficiency. As presented in Figure A4, We investigate changes in parameters and GPU usage of our BISON method. Compared to baseline ER, BISON increases model size by only 0.1% (Split CIFAR-10) and 1.4% (Split CIFAR-100), with GPU usage increasing less than 1%. Since additional classification head is small relative to the backbone, the impact on memory and inference latency is negligible. Overall, our BISON method is computationally efficient and feasible.
Full Implementation Details
Stochastic gradient descent (SDG) optimizer is used for training the parameters from scratch with an initial learning rate $\lambda=$ 0.1 for all experiments except for OCM (OCM/GuoLZ22). Note that the classes are shuffled in each task for every run. As mentioned in the main paper, for our BISON method, we set $\gamma=32$ and $\delta=0.1$ in Eq. (5) for all datasets. For Split CIFAR-100 and Split Mini-ImageNet, $\beta$ and $\lambda$ in Eq. (5) are set to $0.2$ and $10.0$, respectively, while they are set to $0.1$ and $3.0$, respectively, for Split CIFAR-10.
Data Augmentation. We also follow existing works such as (PCR/LinZFLY23) for data augmentation. More specifically, the scale in resized-crop is set to $(0.2,1)$ for each datasets. In color-jitter, the (brightness, contrast, saturation, hue) is set to $(0.4,0.4,0.4,0.1)$ with probability $0.8$. The probability of grey-scale is $0.2$. These data augmentation techniques are applied to both stream and buffer samples for all methods with the exception of OCM, as OCM takes advantage of their own data augmentation techniques.
Reproduction. For ensuring reproducibility, the random seed for all experiments is fixed to be 0, and we use default experimental settings like hyperparameter values for reproducing the results of other compared methods. Unless otherwise notified, we use random sampling to retrieve 10 buffer samples, and employ reservoir updating for buffer management, which are generally adopted by existing replay-based methods. For GSS (GSS/AljundiLGB19), we randomly sample $10$ buffer batches to estimate the maximum cosine similarity scores. For MIR (MIR/AljundiBTCCLP19), the number of sub-sample in their memory retrieval strategy is $50$. For Gdumb (Gdumb/PrabhuTD20), the gradient clip is set to $10$ and we train the memory buffer for $30$ epochs. The data augmentation is applied on their memory training. Memory updating strategy is their greedy balancing updating. For SCR (SCR/MaiLKS21), the temperature $\tau$ in supervised contrastive loss is $0.07$. The projection head is multi-layer perceptron with feature dimensions of $128$. For ASER (ASER/ShimMJSKJ21), the memory updating and retrieval strategy is replaced by their own proposed SV-based strategy. For ER-DVC (DVC/Gu0WD22), the number of sub-sample in their MGI retrieval strategy is $50$, the coefficient of $\lambda_{3}=2$ for Split CIFAR-10, while $\lambda_{3}=4$ for Split CIFAR-100 and Split Mini-ImageNet. We follow the setting $\lambda_{1}=\lambda_{2}=1$ for all datasets. For OCM (OCM/GuoLZ22), we use the reduced ResNet-18, and Adam optimizer is utilized with $0.001$ learning rate, $betas=(0.9,0.99)$ and $0.0001$ weight decay as in their paper. For data augmentation in OCM, resized-crop is set to $(0.3,0.1)$, and the probability of grey-scale is $0.25$. The horizontal-flip, the local rotation and global rotation are also employed. For Split CIFAR-10 and CIFAR-100 in OCM, we follow their hyperparameter values. For Split Mini-ImageNet, we use the same settings as those in Split CIFAR-100. Their max $N_{b}$ is set to 10. For PCR (PCR/LinZFLY23), we use the temperature $\tau=0.07$ in supervised contrastive loss. For LODE (LODE/LiangL23), we take the best hyperparameter values from their paper, where it is equipped with DER++(DERPP/BuzzegaBPAC20) and $\alpha=0.1$, $\beta=1.0$ and $\rho=0.1$ for Split CIFAR-10. As for Split CIFAR-100, $\alpha=0.2$, $\beta=0.2$ and $\rho=0.5$ and we keep the same setting for Split Mini-ImageNet. For all experiments of LODE, the learning rate for optimizing is $0.03$ and the $\beta_{1}=C=1.0$ as default.